OpenStudy (anonymous):

Write the equation that describes each line in slop-intercept form 6. Slope = 5; (1,7) is on the line

OpenStudy (anonymous):

@hockeychick23

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (hockeychick23):

y=5x+2

jimthompson5910 (jim_thompson5910):

y = mx+b y = 5x + b ... plug in m = 5 (the slope) 7 = 5*(1) + b ... plug in the given point (x,y) = (1,7) 7 = 5 + b 7-5 = b 2 = b b = 2 ------------------------------------------------------- m = 5 is the slope b = 2 is the y-intercept ------------------------------------------------------- y = mx+b turns into y = 5x + 2